### v3

# LLMEasyQuant: A User-Friendly Toolkit for LLM Quantization

Abstract
Currently, there are many quantization methods appeared for LLM quantization, yet few are user-friendly and easy to be deployed locally. Packages like TensorRT[NVI24] and Quanto[Fac24] have many underlying structures and self-invoking internal functions, which are not conducive to developers’ personalized development and learning for deployment. Therefore, we develop LLMEasyQuant, it is a package aiming to for easy quantization deployment which is user-friendly and suitable for beginners’ learning.
The code of LLMEasyQuant can be found at code link [^1] https://github.com/NoakLiu/LLMEasyQuant.
1 Introduction
Quantization is the process of mapping a large set of input values to a smaller set of output values, often integers. It is a key technique in digital signal processing where continuous signals are mapped to discrete digital values, and it reduces the data’s precision to make storage and computation more efficient while attempting to retain essential information.
With the development of Large Language Models (LLMs), the models have grown extremely large, so the memory usage and inference speed are greatly limited by the size of the model. Consequently, as one of the most popular technique for model compression, quantization has many variants now used for LLM compression and inference acceleration. The goal of quantization in LLMs is to reduce their size while minimizing its influence on inference speed.
The quantization process can be described by:
$Q(x)=\text{clamp}\left(\left\lfloor\frac{x-\text{min}(X)}{\Delta}+0.5\right\rfloor+z,-128,127\right)$
where:
  • $x$ is a floating-point value from the dataset $X$,
  • $\Delta$ (scale) is calculated as $\frac{\text{max}(X)-\text{min}(X)}{255}$,
  • $z$ (zero point) is $-128$ or calculated to shift the scale,
  • clamp function ensures values are kept within the $[-128,127]$ range.
2 Methodology
LLMEasyQuant is a toolkit designed to simplify the process of deploying quantization on large language models (LLMs). Its functions are included as follows:
  • absmax: $\text{scale}=\frac{127}{\max(|X|)}$
  • zeropoint: Shifting the tensor values based on a computed zero-point
  • symquant: Symmetric quantization, a quantization method for scaling based on the absolute maximum value.[FFBL18]
  • smoothquant: A quantization method to achieve efficient post-training quantizatio for LLMs.[XLS+24]
  • zeroquant: Adjust the numeric range of input data so that zero values in the original data can be represented exactly in the quantized format. [YAZ+22]
  • simquant: A quantization technique by KV Cache Quant. [HKM+24]
3 Implementation
3.1 ZeroQuant: Zero-point Quantization
3.1.1 Definition and Concept
Zero-point quantization, or ZeroQuant [YAZ+22], focuses on adjusting the numeric range of data so that zero values are represented exactly by zero in the quantized format. This technique is particularly useful in optimizing quantization schemes where the preservation of zero values is crucial, such as in sparse datasets used in machine learning and signal processing.
3.1.2 Mathematical Formulation
The quantization process using ZeroQuant can be described by the following formula:
$Q(x)=\text{clamp}\left(\left\lfloor\frac{x-\text{min}(X)}{\Delta}+z\right\rfloor,-128,127\right)$
where:
  • $x$ is the value to be quantized,
  • $\text{min}(X)$ is the minimum value in the dataset,
  • $\Delta$ is the quantization step size, calculated as $\frac{\text{max}(X)-\text{min}(X)}{255}$,
  • $z$ is the zero-point, adjusted so that $\text{min}(X)$ maps to 0,
  • The output is clamped to the range $[-128,127]$, typical of 8-bit signed integers.
3.2 Symmetric 8-bit Quantization
3.2.1 Concept
Symmetric 8-bit quantization [FFBL18] involves mapping both positive and negative values of a dataset uniformly around zero, optimizing the quantization process for symmetric data distributions.
3.2.2 Mathematical Formulation
The symmetric quantization can be described with:
$Q(x)=\text{clamp}\left(\text{round}\left(\frac{x}{s}\right),-128,127\right),\quad s=\frac{\text{max}(|X|)}{127.5}$
where $s$ is the scale factor calculated based on the maximum absolute value of the data, ensuring all quantized values fall within the 8-bit integer range of -128 to 127.
3.3 Layer-by-Layer Quantization
3.3.1 Concept
Layer-by-layer quantization [TOW+23] is a technique applied in deep learning where each layer of a neural network is quantized independently to maintain accuracy while reducing the overall model size and computational demand.
3.3.2 Quantization Procedure
The process involves:
  • Calculating the quantization parameters for each layer separately.
  • Adjusting each layer’s weights and biases according to the quantization rules defined for symmetric or asymmetric quantization methods.
  • Optionally recalibrating layer parameters during or after the quantization to optimize performance.
3.3.3 Implementation Example
Here is an example of applying layer-by-layer quantization:
# Pseudocode for Layer-by-Layer Quantization
for layer in model.layers:
    scale = compute_scale(layer.weights)
    quantized_weights = quantize(layer.weights, scale)
    layer.weights = quantized_weights
3.4 Symmetric 8-bit and Layer-by-Layer Quantization
3.4.1 Layer-by-Layer ZeroQuant Function
Applies quantization per layer in a model:
  • Encode input and compute unquantized outputs.
  • For each layer $i$:
    • Freeze all but layer $i$.
    • Update $i$-th layer’s parameters by quantization.
  • Compute loss and update using:
[…]
  • Optimize with gradient descent.
3.5 SimQuant: A Novel Approach by KV Cache Quant
3.5.1 Definition and Concept
SimQuant [HKM+24], developed by KV Cache Quant, represents a groundbreaking advance in the field of data quantization. This technique is designed to optimize the efficiency of data representation while maintaining high accuracy, particularly in environments where computational resources and storage are limited. Unlike traditional quantization methods that apply fixed parameters across various datasets, SimQuant introduces a dynamic adjustment mechanism that tailors the quantization process to the specific statistical properties of the dataset in use.
3.5.2 Mathematical Formulation
The core of the SimQuant technique is based on the formula:
$Q(x)=\text{round}\left(\frac{x-\min(X)}{\Delta}\right)+Z$
where:
  • $x$ is the data point being quantized,
  • $\min(X)$ is the minimum value in the dataset,
  • $\Delta$ represents the quantization interval, which is adaptively calculated,
  • $Z$ is the zero-point adjustment to center the quantization range.
3.5.3 Algorithmic Steps
  • Data Analysis: Initially, SimQuant analyzes the statistical distribution of the dataset to determine optimal quantization parameters.
  • Parameter Adjustment: It then adjusts $\Delta$ and $Z$ dynamically during the quantization process to minimize information loss.
  • Quantization Application: Finally, the data is quantized using the calculated parameters, ensuring that the most critical information is retained with minimal resource usage.
3.5.4 Quantization Process
  • Calculate range values:
    $\text{vals}_{\text{min}}=\min(X_{\text{channel}}),\quad\text{vals}_{\text{max}}=\max(X_{\text{channel}})$
  • Compute scale $s$ and zero point $z$:
    $s=\frac{2^{\text{bits}}-1}{\text{vals}_{\text{max}}-\text{vals}_{\text{min}}},\quad z=-\text{vals}_{\text{min}}\cdot s$
  • Apply quantization:
    $X_{\text{quant}}=\text{clamp}(\lfloor X\cdot s+z+0.5\rfloor,0,2^{\text{bits}}-1)$
3.5.5 Dequantization Process
$X_{\text{dequant}}=\frac{X_{\text{quant}}-z}{s}$
3.5.6 Pseudocode for SimQuant
Algorithm 1 KVQuant Quantization
1:function QuantizeKVCache(Key, Value, scaling_factor, zero_point)
2:     for each layer $l$ do
3:         $Key_{l}\leftarrow\text{Quantize}(Key_{l},scaling\_factor,zero\_point)$
4:         $Value_{l}\leftarrow\text{Quantize}(Value_{l},scaling\_factor,zero\_point)$
5:     end for
6:     return $Key,Value$
7:end function
3.6 SmoothQuant: Smoothing Quantization Process
3.6.1 Definition and Concept
SmoothQuant [XLS+24] is a post-training quantization (PTQ) method designed to facilitate accurate and efficient quantization of large language models (LLMs). It aims to achieve 8-bit quantization for both weights and activations without retraining, ensuring minimal loss in model accuracy. SmoothQuant addresses the challenge of activation quantization by redistributing the quantization difficulty from activations to weights, which are inherently easier to quantize.
The key concept behind SmoothQuant is the smoothing of activation outliers, which typically hinder effective quantization. By migrating the quantization difficulty from activations to weights, SmoothQuant ensures that both can be quantized to 8-bit integers (INT8) effectively. This migration is achieved through a mathematically equivalent transformation that adjusts the scale of activations and weights.
3.6.2 Mathematical Formulation
The quantization and smoothing process of the ‘SmoothQuantMatrix‘ can be mathematically described as follows:
[…]
where the scale factor $s$ is calculated using the activity scales act_scales and the weight scales of the features:
$s=\left(\frac{\text{act\_scales}^{\alpha}}{\text{weight\_scales}^{1-\alpha}}\right)$
Here, $\alpha$ is a parameter that determines the balance between the activity and the weight scales, affecting the smoothness of the quantization.
3.6.3 SmoothQuant Methodology
SmoothQuant smooths activation outliers by migrating quantization difficulty from activations to weights. The following pseudocode outlines the main steps of the SmoothQuant process.
Algorithm 2 SmoothQuant Process
1:Input: Activation tensor $X$, Weight tensor $W$
2:Output: Quantized activation $\overline{X}$, Quantized weight $\overline{W}$
3:Compute per-channel scaling factor $s_{j}=\max(|X_{j}|)^{\alpha}/\max(|W_{j}|)^{1-\alpha}$
4:for each channel $j$ do
5:     Smooth activation $X_{j}\leftarrow X_{j}/s_{j}$
6:     Adjust weights $W_{j}\leftarrow W_{j}\times s_{j}$
7:end for
8:Quantize smoothed activation and adjusted weights
[…]
11:return $\overline{X},\overline{W}$
3.6.4 Implementation Details
The ‘SmoothQuantMatrix‘ class computes the smoothing scales based on the provided activity scales and the inherent data characteristics, dynamically adjusting each feature’s scale. This customization allows the quantization process to be more adaptive and sensitive to the underlying data distribution.
4 result
Figure 1: Quantized Weights Distribution
Figure 2: Performance Comparison after Quantization on GPT
Models Perplexity (ppl)
GPT-2 4.01
GPT-2 INT8 6.83
GPT-2 AbsMax Quantize 9.32
GPT-2 ZeroPoint Quantize 8.93
GPT-2 Smooth Quant Apply 6.31
GPT-2 Sim Quantize 7.16
GPT-2 Sym Quantize 8bit 7.01
GPT-2 Sym Quantize 8bit ZeroQuant Func 7.37
Table 1: Perplexity Analysis of Quantization Models
5 Discussions
LLMEasyQuant combined with downstream applications such as [WLC+24], [LZC24], [LWC+24] can achieve satisfactory performance.
5.1 Simplification of Quantization Processes
LLMEasyQuant provides a user-friendly interface that simplifies the application of quantization techniques, making it accessible to both novices and experienced users without requiring deep technical knowledge of the underlying algorithms.
5.2 Customization and Flexibility
Despite its simplicity, the package offers extensive customization options that allow users to tailor the quantization process to their specific needs, balancing efficiency and performance according to the model’s deployment context.
5.3 Efficiency in Deployment
Optimized for performance, LLMEasyQuant helps reduce the computational load and memory usage of models, facilitating their deployment on devices with limited resources such as mobile phones and embedded systems.